OpenStudy (anonymous):

Convert the mixed number, 1 3/4 , to a fraction. Reduce fraction if necessary

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Fractions of the form \(a\frac{b}{c}\) represent: \[a\frac{b}{c}=a+\frac{b}{c}=\frac{ac+b}{c}\] So in this case, we know our a, b, and c. And we can plug it in and simplify: \[1\frac{3}{4}=\frac{(1)(4)+(3)}{(4)}\] I will leave its computation to you!